OpenStudy (anonymous):

what is the solution of the matrix equation? [ 5 -2] X= [ 2 ] [ 2 -1] [-4]

OpenStudy (unklerhaukus):

\[\begin{bmatrix}5&-2\\2&-1\end{bmatrix}\begin{bmatrix}x_1\\x_2\end{bmatrix}=\begin{bmatrix}2\\-4\end{bmatrix}\] \[\left[\begin{array}{cc|c}{}5&-2&2\\2&-1&-4\end{array}\right]\]

OpenStudy (unklerhaukus):

perform row operations until you get \[\left[\begin{array}{cc|c}{}1&0&x_1\\0&1&x_2\end{array}\right]\]

OpenStudy (unklerhaukus):

change the first row, first \[R_1\to R_1-2R_2\]